OpenStudy (anonymous):

A swimming pool can be filled in 18 hours if water enters through a pipe alone, or in 25 hours if water enters through a hose alone. If water is entering through both the pipe and the hose, how long will it take to fill the pool 5/6 full?

OpenStudy (anonymous):

P = pipe flow rate H = hose flow rate t = time P = 1 pool / 18 hours = 1/18 H = 1 pool / 25 hours = 1/25 Pt + Ht = 5/6 t(P + H) = 5/6 t(1/18 + 1/25) = 5/6 t(25/450 + 18/450) = 5/6 t(43/450) = 5/6 t = (450*5)/(43*6) = approximately 8.72093023

OpenStudy (anonymous):

Ask me any questions you have about what I did or why.

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

Yes Thank you! =)

OpenStudy (anonymous):

You're welcome =D